Shift report 20010805 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 05 2001 - 08:09:00 EDT

  • Next message: Apache: "Shift report 20010805 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Aug  5 08:08:59 2001, Shift: 00:00-08:00
    Supervisor: KH, with: A. Hanssen
    Log book#7 pages 20 - 23
    
    Shift Summary:
    ------------------------------------------------------------
    Two stores with two runs
    1:50 - 2:10 run 3993  ~43936 collisions
    Then yellow abort kicker tripped, pulling the quench link
    4:15 - 8:00 run 3994  still running at shift end
                          estimate 200-300k collisions but
                          it could be much more.  Based on
                          SuperMon which was not running the
                          whole time.
    DAQ still taking data at 40 events/sec at shift end
    
    Running T1 anode at 1100V
    We see tracks in T1 some of which seem to match up with
    tracks in T2.  A few pictures are in the log book.
    
    No known detector problems.
    
    Machine had only problems mentioned above.
    One unsuccessful store due to a chipmunk put beam stops in
    and caused the ramp to fail.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 05 2001 - 08:09:50 EDT